OpenStudy (anonymous):

The length of a rectangle is three times its width. If the perimeter of the rectangle is 56in , find its area.

OpenStudy (anonymous):

if width is x,what's its length?

OpenStudy (anonymous):

Hint: Formula for perimeter is 2(w)+2(l).

OpenStudy (anonymous):

where w is width and l is length.

OpenStudy (anonymous):

Length= 3 x width Let width= w Perimeter = 2(width)+2(length) 56=2(w)+2(3w) 56=2w+6w Solve for w. After finding w, you can find the length by substitution w=? into the length equation. After finding the length, you can now find the area which formula is... Length x Width